OpenStudy (anonymous):

Express y = 14 - 3x - x^2 in the form y = p(x + q)^2. Hence, find the maximum value of y and the value of x at which the maximum value occurs.

sam (.sam.):

\[\frac{65}{4}-\left(x+\frac{3}{2}\right)^2\]

sam (.sam.):

Maximum at vertex (-3/2, 65/4) Because the coefficient of x^2 is negative, its a maximum

OpenStudy (anonymous):

Thanks again! Any idea how to sketch the graph for y = 14 - 3x - x^2?

OpenStudy (anonymous):

THANK YOU, Sam! (:

sam (.sam.):

welcome :)